In a recent​ year, companies spent a total of ​$87.6 billion on​ newspaper, television, and radio ads. The total amount spent on television and radio ads was only ​$2.8 billion more than the amount spent on newspaper ads alone. The amount spent on newspaper ads was​ $5.5 billion more than what was spent on television ads. How much was spent on each form of​ advertising? ​(​Hint: Let the variables represent numbers of billions of​ dollars.)
Answer by stanbon(75887)

In a recent​ year, companies spent a total of ​$87.6 billion on​ newspaper, television, and radio ads. The total amount spent on television and radio ads was only ​$2.8 billion more than the amount spent on newspaper ads alone. The amount spent on newspaper ads was​ $5.5 billion more than what was spent on television ads. How much was spent on each form of​ advertising? ​(​Hint: Let the variables represent numbers of billions of​ dollars.)
Equations:
n + t + r = 87.6
t + r = n + 2.8
n = t + 5.5
------------------
Rearrange::
n + t + r = 87.6
n - t - r = -2.8
n - t + 0 = 5.5
----------------------
Use any method you know to solve the system to get:
n = 42.4
t = 36.9
r = 8.3
---
I used my TI-84+ Matrix function.
Cheers,
Stan H.